OpenStudy (yonce):

Which three lengths could be the lengths of the sides of a triangle? 11 cm, 6 cm, 17 cm 6 cm, 22 cm, 10 cm 8 cm, 15 cm, 22 cm 22 cm, 5 cm, 10 cm

jimthompson5910 (jim_thompson5910):

Rule: if a triangle has sides p,q,r, then all of the following must be true p+q > r p+r > q q+r > p if any of the above inequalities are false, then a triangle can't be formed

jimthompson5910 (jim_thompson5910):

basically, if you can add up any two sides and the result is NOT larger than the third side, then a triangle is NOT possible for example, choice A has 11 cm, 6 cm, 17 cm 11+6 = 17 but the result of the addition(17) is NOT larger than the other side (17) so a triangle is NOT possible for choice A

jimthompson5910 (jim_thompson5910):

choice B is the same 10+6 = 16 but it is not larger than 22 so choice B is also not possible

OpenStudy (yonce):

So would it be C?

jimthompson5910 (jim_thompson5910):

yes because D is out too 5+10 = 15 but it's not larger than 22

jimthompson5910 (jim_thompson5910):

8 cm, 15 cm, 22 cm 8+15 = 23 which is larger than 22 8+22 = 30 which is larger than 15 15+22 = 37 which is larger than 8 all three inequalities (shown above) hold true, so a triangle is possible

OpenStudy (yonce):

Thanks

jimthompson5910 (jim_thompson5910):

np